OpenStudy (anonymous):

Ok I need help! I am doing an experiment and don't understand what they want me to do. I have the masses of everything, but I don't know if the mole calculations are right. I don't know how to get the chemical formula either. Mass of CuSO4 xH2O: 5.000g Mass of CuSO4: 3.196g Mass of H2O: 1.804g Moles of CuSO4: 0.02 Moles of H2O:0.10 Moles of H2O for every mole of CuSO4: Chemical formula of CuSO4 XH2O (Hint: replace the X with the number of moles of water for every mole of CuSO4):

OpenStudy (anonymous):

I will medal and fan!

OpenStudy (anonymous):

@satellite73 @nincompoop

OpenStudy (photon336):

Because like if you take the molar mass of CUSO4 and then convert, the grams you have to moles this is what you get Mass of CuSO4: 3.196g \[\frac{ mol }{ 346grams }*3.196g = 0.009 moles, CUSO_{4}\] \[\frac{ mol }{ 18grams }*(1.84g) = 0.10 moles\] H2O

OpenStudy (photon336):

We take the molar ratio of the two dividing # of moles of water by # of moles of CUSO4 I rounded to the nearest whole number. \[\frac{ H_{2}O }{ CUSO_{4} } = \frac{ .10 }{ .01 } = 10 moles\] so we have 10 moles of water per every mole of CUSO4 so it would be CUSO4*10H2O

OpenStudy (photon336):

@LiloLaurna How did you get 0.02 moles of CUSO4?

OpenStudy (anonymous):

I used an online calculator but it didn't show me the work. Thank you @Photon336 I had no clue how to do this and you helped a lot!

OpenStudy (rushwr):

@Photon336 But I think the moles of CuSO4 u have taken is wrong Moles = mass divided by molar mass and the molar mass of CuSO4 is 159.609gmol^-1 ( So if we round it off tht would be 160) \[Moles = \frac{ 3.196g }{ 160gmol ^{-1} }= 0.02 moles\] \[\frac{ Moles of H _{2}O }{Moles of CuSO _{4} } = \frac{ 0.1 }{ 0.02 } = 5 \] So I guess lilo is correct ! SO for every mole of CuSO4 u find 5 moles of water. (I think i didn't do any other mistakes here )

OpenStudy (photon336):

Thanks a lot @rushwr strange I looked it up twice.

OpenStudy (rushwr):

ya no problem ! Didn't notice that u were wrong! I was wondering how lilo's calculator went wrong ! @Photon336